OpenStudy (anonymous):

hi, i'm having trouble with the integral of x/(2y+1)dy

OpenStudy (anonymous):

I need to know the relationship between x and y for that to be done. Did you perhaps mean the following?\[\int \frac{y}{2y+1}dy\]

OpenStudy (anonymous):

(Unless you intended x as a constant?)

OpenStudy (anonymous):

nope, its a part of a larger question. i'm finding the double integral of (x^2 + x/(2y+1)). so to start it off i need to integrate the whole thing with respect to y so it would be (yx^2) for the first term, but i don't know how to do the second term (original question)

OpenStudy (anonymous):

but, basically x is a constant to start with

OpenStudy (anonymous):

use substitution, the integral of 1/u is ln u

OpenStudy (anonymous):

alrighty, i was afraid i'd have to do that..

OpenStudy (anonymous):

du is 2dx so you need to get a 2 there

OpenStudy (anonymous):

wouldn't it be 2dy?

OpenStudy (anonymous):

yes that's what i meant

OpenStudy (anonymous):

HAHAHA "alrighty, i was afraid i'd have to do that.."

OpenStudy (anonymous):

don't know if it'll work cuz of that first x^2 term..

OpenStudy (anonymous):

you integrate the two terms seperately

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

so the integral of x^2 is just x^2*y

OpenStudy (anonymous):

\[\int {\frac{x}{{2y + 1}}dy = \frac{1}{2}x\int {\frac{2}{{2y + 1}}} } dy = \frac{1}{2}x\int {\frac{1}{u}} du = \frac{1}{2}x\ln u = \frac{1}{2}x\ln 2y + 1\]

OpenStudy (anonymous):

thanks